# QMetro++ - Python optimization package for large scale quantum metrology with customized strategy structures

### v1

@@ Appendix A Choi-Jamiołkowski matrix and the link product
When discussing networks of channels it is useful to use a notion of the Choi-Jamiołkowski matrix (CJ). For a linear map
@@
its CJ matrix is given by [36, 38]:
$P=\left(\mathrm{P}\otimes\mathbb{I}_{\mathcal{A}}\right)\left(\ket{\Phi}\bra{\Phi}\right)\in\mathcal{L}\left(\mathcal{B}\otimes\mathcal{A}\right),$ (56)
@@
    if $\mathrm{P}=\mathrm{Tr}_{\mathcal{A}}$ represents tracing out over the space $\mathcal{A}$ than $P$ is the identity matrix on $\mathcal{A}$,
@@
[…]
@@
[…]
@@
[…]
@@
[…]
where $\mathrm{Tr}_{Z}$, $\text{1}_{Z}$ and $T_{Z}$ are trace, identity matrix and transposition for $\bigotimes_{i\in Z}\mathcal{H}_{i}$ and it is assumed that there are SWAP operators put inside the trace such that terms in round brackets act on the tensor product of spaces $\mathcal{H}_{i}$ in the same order. Additionally, since the relation between channels and CJ matrices is isomorphic [36, 38] link product can be naturally extended to channels.
@@ Appendix B Tensor network formalism and tensor representation
It is convenient to denote tensors and their contractions using diagrams. In this formalism tensor $T^{\mu_{0}\mu_{1}...\mu_{n-1}}$ is represented as a vertex of a graph with edge for each index:
@@
Tensor network formalism can be used to conveniently express link product. For a Hilbert space $\mathcal{H}=\bigotimes_{i=0}^{n-1}\mathcal{H}_{i}$ with $d_{i}:=\dim\mathcal{H}_{i}$ and a linear operator acting on it:
$M=\sum_{\begin{subarray}{c}a_{0},...,a_{n-1}\\ b_{0},...,b_{n-1}\end{subarray}}M^{a_{0}...a_{n-1}}_{b_{0}...b_{n-1}}\ket{a_{0}...a_{n-1}}\bra{b_{0}...b_{n-1}},$ (64)
we define tensor representation of $M$ to be a tensor:
@@
that is only physical indices of $M_{i}$ are joined. This combined with the use of ‘*’ to denote both link product and tensor contraction allows for convenient depiction of link product with CJ matrices which are MPOs.
@@
Completely Positive Trace Preserving

### v2

@@ Appendix A Choi-Jamiołkowski matrix and the link product
When discussing networks of channels it is useful to introduce the notion of the Choi-Jamiołkowski matrix (CJ). For a linear map
@@
the corresponding CJ matrix is given by [12, 5]:
[…]
@@
    if $\mathrm{P}=\mathrm{Tr}_{\mathcal{A}}$ represents the partial trace with respect to $\mathcal{A}$ than $P$ is the identity matrix on $\mathcal{A}$,
@@
    $\mathrm{P}$ is trace-preserving if and only if $\mathrm{Tr}_{\mathcal{B}}P=\text{1}_{\mathcal{A}}$,
@@
    $\mathrm{P}$ is unital, that is $\mathrm{P}(\text{1}_{\mathcal{A}})=\text{1}_{\mathcal{B}}$, if and only if $\mathrm{Tr}_{\mathcal{A}}P=\text{1}_{\mathcal{B}}$.
@@
$R=P*Q=\mathrm{Tr}_{2}\left[\left(\text{1}_{34}\otimes P^{T_{2}}\right)\left(Q\otimes\text{1}_{01}\right)\right],$ (59)
@@
$=\mathrm{Tr}_{X\cap Y}\left[\left(\text{1}_{Y\setminus X}\otimes M^{T_{X\cap Y}}\right)\left(N\otimes\text{1}_{X\setminus Y}\right)\right],$
where $\mathrm{Tr}_{Z}$, $\text{1}_{Z}$ and $T_{Z}$ are trace, identity matrix and transposition for $\bigotimes_{i\in Z}\mathcal{H}_{i}$ and it is assumed that there are SWAP operators inserted inside the trace such that terms in round brackets act on the tensor product of spaces $\mathcal{H}_{i}$ in the same order. Additionally, since the relation between channels and CJ matrices is isomorphic [12, 5] link product can be naturally extended to channels.
@@ Appendix B Tensor network formalism and tensor representation
It is convenient to denote tensors and their contractions using diagrams. In this formalism tensor $T^{\mu_{0}\mu_{1}...\mu_{n-1}}$ is represented as a vertex of a graph with an edge for each index:
@@
Tensor network formalism can be used to conveniently express the link product. For a Hilbert space $\mathcal{H}=\bigotimes_{i=0}^{n-1}\mathcal{H}_{i}$ with $d_{i}:=\dim\mathcal{H}_{i}$ and a linear operator acting on it:
$|,M=\sum_{\begin{subarray}{c}a_{0},...,a_{n-1}\\ b_{0},...,b_{n-1}\end{subarray}}M^{a_{0}...a_{n-1}}_{b_{0}...b_{n-1}}\ket{a_{0}...a_{n-1}}\bra{b_{0}...b_{n-1}}$, (64)
we define the tensor representation of $M$ to be a tensor:
@@
that is, only physical indices of $M_{i}$ are joined. This, combined with the use of ‘*’ to denote both link product and tensor contraction, allows for convenient depiction of link product with CJ matrices which are MPOs.
@@
Completely Positive Trace-Preserving